Power Drive Corporation designs and produces a line of golf equipment and golf apparel. Power Drive has 100,000 shares of common stock outstanding as of the beginning of 2018. Power Drive has the following transactions affecting stockholders' equity in 2018.
  

March 1 Issues 61,000 additional shares of $1 par value common stock for $58 per share.

May 10 Repurchases 5,600 shares of treasury stock for $61 per share.

June 1 Declares a cash dividend of $1.80 per share to all stockholders of record on June 15. (Hint:  Dividends are not paid on treasury stock.)

July 1 Pays the cash dividend declared on June 1.

October 21 Reissues 2,800 shares of treasury stock purchased on May 10 for $66 per share.
  

Power Drive Corporation has the following beginning balances in its stockholders' equity accounts on January 1, 2018: Common Stock, $100,000; Additional Paid-in Capital, $5,100,000; and Retained Earnings, $2,600,000. Net income for the year ended December 31, 2018, is $660,000.
  

Required:

Prepare the statement of stockholders’ equity for Power Drive Corporation for the year ended December 31, 2018. (Amounts to be deducted should be indicated by a minus sign.)

POWER DRIVE CORPORATION Statement of Stockholders Equity For the Year Ended December 31, 2018 Common Stock Paid-in Capital APlease explain how you got every number.. thank you.
Answer #1

Statement of stockholder's equity

For the year ended December 31, 2018

Common Stock Additional paid in capital Retained earnings Treasury Stock total Stockholder's equity
Balance, January 1 100000 5100000 2600000 7800000
Issue common Stock 61000 61000*57 = 3477000 3538000
Purchase treasury stock 5600*61 = -341600 -341600
Dividends -279720 -279720
Sale of treasury stock 2800*5 = 14000 2800*61 = 170800 184800
Net income 660000 660000
Balance December 31 161000 8591000 2980280 -170800 11561480